OpenStudy (anonymous):

The graph of y=|x| is reflected in the x-axis and then moved 1 unit to the left and 3 units up. Write an equation of the new function.

OpenStudy (solomonzelman):

okay, here are some rules, SHIFTS from \(\large\color{black}{ f(x) }\) to form \(\large\color{black}{ g(x) }\). \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~g(x)=\left| x \color{blue}{ -~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm{c} }\) units to the right. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~g(x)=\left| x \color{blue}{ +~\rm{c} }\right| }\) \(\large\color{blue}{ ~\rm{c} }\) units to the left. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~g(x)=\left| x \right| \color{blue}{ +~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units to up. \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~g(x)=\left| x \right| \color{blue}{ -~\rm{c} }}\) \(\large\color{blue}{ ~\rm{c} }\) units to down.

OpenStudy (anonymous):

when you reflect over the x axis u are ultimately making your y negative

OpenStudy (solomonzelman):

the reflection across x-axis, \(\large\color{black}{ f(x)=\left| x \right| ~~~~~\bf{\rightarrow}~~~~~g(x)=\color{blue}{ ~\rm{-} }\left| x \right| }\)

OpenStudy (solomonzelman):

So you get it?

OpenStudy (anonymous):

Yes thank you!!

OpenStudy (solomonzelman):

can you write or draw your new equation (just in case, you don't have to) ?

OpenStudy (anonymous):

\[y=-\left| x+1 \right|+3\]

OpenStudy (solomonzelman):

YES !!

OpenStudy (anonymous):

thank you!!

OpenStudy (solomonzelman):

beautiful, and with latex, \(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ \large\color{black}{ y=\color{red}{ ~\rm{- }} \left| x \color{blue}{ ~\rm{+~1} }\right|\color{green}{ ~\rm{+~3} }} }}}\)

OpenStudy (solomonzelman):

like this, \(\huge\color{ blue }{\huge {\bbox[5pt, lightcyan ,border:2px solid red ]{ \large\color{black}{ y=\color{red}{ ~\rm{- }} \left| x \color{blue}{ ~\rm{+~1} }\right|\color{green}{ ~\rm{+~3} }} }}}\)

OpenStudy (solomonzelman):

Anyways..... good luck with your math:) You welcome!

OpenStudy (anonymous):

Like the boxes!